OpenStudy (anonymous):

What is the relationship between the lines determined by the following two equations? 6x – 2y = 16 y = 3x – 8 parallel perpendicular neither parallel nor perpendicular they are the same line

OpenStudy (lncognlto):

First, let's take 6x – 2y = 16 and simplify it. Therefore -2y = -6x + 16 Therefore y = 3x -8 Which is the same as the second equation given. Therefore the lines are the same.

OpenStudy (anonymous):

Can you help me with a few more? c:

OpenStudy (anonymous):

Which of the following has a slope of –2 and a y-intercept of 4? y = –2x + 4 y = 2x + 4 y = 2x – 4 y = –2x – 4

OpenStudy (lncognlto):

Sure, I can try. ^-^

OpenStudy (lncognlto):

In the standard equation y = mx + c, the value of m is the slope of the line, and the value of c is the y-intercept. So, comparing the given equations to this standard form, the slopes and y-intercepts are as follows: -2 and 4, 2 and 4, 2 and -4, -2 and -4. Therefore the first equation is the one we are looking for.

OpenStudy (anonymous):

these are of the form: y=mx+c where m=slope and c=y intercept now try to do tis on ur own

OpenStudy (anonymous):

Ah that makes sense! thank you, I have I think 2 more if you dont mind helping me, I really really appreciate it! @inco

OpenStudy (lncognlto):

Ok, sure!

OpenStudy (anonymous):

The graph of a line has a slope of 3 and passes through the point (–4, –5). What is the equation of the line in point-slope form? y + 4 = 3(x + 5) y – 4 = 3(x – 5) y + 5 = 3(x + 4) y – 5 = 3(x – 4)

OpenStudy (lncognlto):

Well, I'm not exactly sure how this one is meant to be solved, but this is how I'd do it: To find the equation of the line required, we suubstitute m = 3 and (-4, -5) into y = mx +c. Therefore, -5 = 3(-4) + c Therefore, -5 = -12 + c Therefore, c = 7 So the equation of the line equals y = 3x + 7 Now we simply the options given: y + 4 = 3(x + 5) -> y = 3x + 11 y – 4 = 3(x – 5) -> y = 3x - 11 y + 5 = 3(x + 4) -> y = 3x + 7 y – 5 = 3(x – 4) -> y = 3x - 7 Therefore, the required equation is the third one. On a side note, the way I just solved it is probably an unnecessarily long way of doing so, but I cannot think of another way right now.

OpenStudy (anonymous):

Okay, that makes sense, I'll write this down. Two more then I think I'm done! I really appreciate your time and help. What are the x- and y-intercepts of the line? 2x – 3y = –12 x-intercept 6, y-intercept –4 x-intercept –6, y-intercept 4 x-intercept –4, y-intercept 6 x-intercept 4, y-intercept –6 Question 12.12. Select the equation of the line that passes through the point (–2, –1) and has slope 5 in point-slope form. (y – 1) = 5(x – 2) (x + 2) = 5(y – 1) (x – 2) = 5(y + 1) (y + 1) = 5(x + 2)

OpenStudy (lncognlto):

Right. To find the intercepts of the line 2x – 3y = –12, we proceed as follows. For the y-intercept, when x = 0, 2(0) - 3y = -12. Therefore, -3y = -12 Therefore, y = 4 For the x-intercept, when y = 0, 2x - 3(0) = -12 Therefore, 2x = -12 Therefore, x = -6 Therefore the y-intercept is 4, and the x-intercept is -6. In general, to find the y-intercept, we substitute x = 0 into the equation, and to find the x-intercept, when substitute in y = 0.

OpenStudy (anonymous):

Ah! Thank you! c:

OpenStudy (anonymous):

I don't quite understand the last one?

OpenStudy (lncognlto):

For the last one, I think I'm going to solve it like I did that earlier one I wasn't sure about. So here goes: To find the equation of the line required, we suubstitute m = 5 and (-2, -1) into y = mx +c. Therefore, -1 = 5(-2) + c Therefore, -1 = -10 + c Therefore, c = 9 So the equation of the line equals y = 5x + 9 Now we simply the options given: (y – 1) = 5(x – 2) -> y - 1 = 5x - 10 -> y = 5x - 9 (x + 2) = 5(y – 1) -> x + 2 = 5y - 5 -> 5y = x + 7 -> y = (1/5)x + 7/5 (x – 2) = 5(y + 1) -> x - 2 = 5y + 5 (This isn't the one either. XD) (y + 1) = 5(x + 2) -> y + 1 = 5x + 10 -> y = 5x + 9 Therefore, the required equation is the last one.

OpenStudy (anonymous):

Ohhh! okay, thank you so much for taking your time to help me c:

OpenStudy (lncognlto):

You're welcome! :D